From one hospital, a small family rides their car and travels 25 km in a direction 30° south of west and then 30 km toward the north to a second hospital in search of a vacancy to accommodate a severe-case CoViD-19 relative patient. Where is the second hospital with respect to the first hospital, i.e. direction of the resultant displacement?


1
Expert's answer
2021-10-26T10:47:24-0400


Take a component of x-axis and y-axis

Resultant (R)=25cos30°km+30 km

R=21.65+30=51.65Km

R=51.65 Km

Now direction

"\phi=tan^{-1}(\frac{25}{30})=39.80°"

Direction 39.80 ° N to W

N 39.80°W